OpenStudy (anonymous):

Which ratio is the same proportion as 18 : 48? A.3/8 B.1/30 C.6/8 D.1/3

OpenStudy (lizz123):

You divide by 6 each side wht is your answer

OpenStudy (lizz123):

18/6=3 48/6=?

OpenStudy (acxbox22):

simplify the ratio like you would to a fraction

OpenStudy (anonymous):

48/6=8

OpenStudy (lizz123):

\[48\div6\]

OpenStudy (lizz123):

so since 18/6=3 and 48/6=8 you are left with 3:8

OpenStudy (anonymous):

thatis not one of the answers

OpenStudy (lizz123):

3/8

OpenStudy (acxbox22):